# Combinatorial Selection with Costly Information

### v1

@@ E.3 Amortization for MDPs
and by defining $\hat{Z}_{j}:=\min\{g_{j},Z_{j}\}$, we have
which in turn allows us to argue that
for all $y\in\real$ and $j\in[k]$.
  1. 1.
  2. 2.
We can now define the amortized cost function $\rho$ and the cost sharing vector $b$. Fix some $t\in T$, and let $j\in[k]$ and $s\in S$ be the indices of the unique action $a_{j}$ and state $s\in R_{j}$ such that $t\in T(s)$. We define
$\rho(t):=m_{j}(\min\{g_{j},\rho_{s}(t)\})$
and
$b_{\sigma t}:=\operatorname{E}\left[\rho_{s}(t)\right]-\operatorname{E}\left[\rho(t)\right]$
and for all other $s^{\prime}\in S\setminus\{\sigma\}$ with $t\in T(s^{\prime})$, we use the same cost share $b_{s^{\prime}t}=b^{s}_{s^{\prime}t}$ that was used in $\mathcal{M}_{s}$. The proof of the three properties follows precisely the same steps as Lemma 3.3. ∎
From Lemma E.3, we immediately obtain the following counterpart of Theorem 3.4; the proof follows exactly the same steps as the proof of Theorem 3.4 that we presented in Appendix A, and is thus omitted.
Theorem E.4.
For any matroid-max-CICS instance $\mathbb{I}=(\mathcal{M}_{1},\cdots,\mathcal{M}_{n},\mathcal{F})$, the expected utility of the optimal adaptive policy is at most $\operatorname{E}\left[\max_{S\in\mathcal{F}}\sum_{i\in S}W^{*}_{\mathcal{M}_{i}}\right]$.
E.4 Local Approximation
Definition 15 (Local approximation for maximization settings).
Let $\mathcal{M}$ be any MDP. We say that a committing policy $\pi\in\mathcal{C}\left(\mathcal{M}\right)$ $\alpha$-locally approximates $\mathcal{M}$ for some $\alpha\in(0,1]$ if
$f_{\mathcal{M}^{\pi}}(\alpha y)\geq\alpha\cdot f_{\mathcal{M}}(y)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall y\in\real.$
Theorem E.5 (Composition theorem for maximization settings).
Let $\mathbb{I}=(\mathcal{M}_{1},\cdots,\mathcal{M}_{n},\mathcal{F})$ be any instance of matroid-max-CICS, where each MDP $\mathcal{M}_{i}$ admits an $\alpha$-local approximation under some committing policy $\pi_{i}\in\mathcal{C}\left(\mathcal{M}_{i}\right)$. Then,
$\operatorname{E}\left[\max_{S\in\mathcal{F}}\sum_{i\in S}W^{*}_{\mathcal{M}_{i}^{\pi_{i}}}\right]\geq\alpha\cdot\operatorname{E}\left[\max_{S\in\mathcal{F}}\sum_{i\in S}W^{*}_{\mathcal{M}_{i}}\right].$
By combining Theorem E.5 with the optimality of the Water Draining Index policy (Theorem E.2) and the upper bound of Theorem E.4, we obtain a way to efficiently approximate the optimal solution for matroid-max-CICS assuming that the underlying MDPs achieve good local approximation guarantees.
Appendix F The Combinatorial Setting
In this section, we describe how our entire framework extends to combinatorial settings beyond the case of matroids. We will state our results in full generality and distinguish between minimization and maximization whenever needed. The framework we consider is based on [Singla, 2017] and its followup [Gupta et al., 2019], where the authors develop a technique for combinatorial selection over Markov chains. In this section, we show that under local approximation, their results can be seamlessly extended to arbitrary MDPs.
Definition 16 (CICS).
$\sum_{t\in T}v(t)+h(S)+\sum_{a\in A}c(a)$
$\sum_{t\in T}v(t)+h(S)-\sum_{a\in A}c(a).$
Theorem F.1.
Let $\mathbb{I}=(\mathcal{M}_{1},\cdots,\mathcal{M}_{n},\mathcal{F},h)$ be any instance of CICS. For each $i\in[n]$, let $W^{*}_{\mathcal{M}_{i}}$ be the water filling (resp. water draining) surrogate costs in the minimization (resp. maximization) setting. Let $\mathrm{OPT}$ be the expected cost (resp. utility) of the optimal adaptive policy. Then:
  1. 1.
    For the minimization setting, $\mathrm{OPT}\geq\operatorname{E}\left[\min_{S\in\mathcal{F}}(\sum_{i\in S}W^{*}_{\mathcal{M}_{i}}+h(S))\right]$.
  2. 2.
    For the maximization setting, $\mathrm{OPT}\leq\operatorname{E}\left[\max_{S\in\mathcal{F}}(\sum_{i\in S}W^{*}_{\mathcal{M}_{i}}+h(S))\right]$.
Theorem F.2.
Let $\mathbb{I}=(\mathcal{M}_{1},\cdots,\mathcal{M}_{n},\mathcal{F},h)$ be any instance of CICS. For each $i\in[n]$, let $W^{*}_{\mathcal{M}_{i}}$ be the water filling (resp. water draining) surrogate costs in the minimization (resp. maximization) setting. Finally, for each $i\in[n]$, let $\pi_{i}\in\mathcal{C}\left(\mathcal{M}_{i}\right)$ be some committing policy that $\alpha$-locally approximates $\mathcal{M}_{i}$. Then:
  1. 1.
  2. 2.
By combining Theorem F.1 with Theorem F.2, and assuming that our MDPs admit local approximation guarantees, we are left with the task of optimizing over the CICS instance that is generated by the committing policies; observe that this is now an instance over Markov chains. In the case of matroid constraints, we showed that we can always efficiently achieve this via the water filling/draining index policy. However, depending on the combinatorial constraint, efficient optimization might not be possible – consider for example the case where each MDP is a single terminal state corresponding to a set of elements and we need to select a minimum cost set cover.
The final key to the puzzle will be a way of efficiently approximating the optimal policy for a CICS instance $\mathbb{I}=(\mathcal{M}_{1},\cdots,\mathcal{M}_{n},\mathcal{F},h)$ over Markov chains. This is precisely the setting that is considered by Gupta et al. [2019]. In this work, it is shown that a sufficient condition to get an efficient $\beta$-approximation to the optimal policy for a CICS instance over Markov chains is for the underlying pair $(\mathcal{F},h)$ to admit a $\beta$-approximate frugal algorithm. We defer the reader to Gupta et al. [2019] for the full details. Here, we will just mention some examples of such combinatorial settings:
  • Matching constraints in the maximization setting admit a $2$-approximate frugal algorithm.
Corollary F.3.
  1. 1.
    Each MDP $\mathcal{M}_{i}$ admits an $\alpha$-local approximation.
  2. 2.
Then, we can efficiently obtain an $(\alpha\cdot\beta)$-approximation to the optimal adaptive policy for $\mathbb{I}$.
We are left with the task of proving our extended Theorems F.1 and F.2. It is not hard to see that both proofs are identical to their matroid counterparts and are thus omitted. In particular:
  1. 1.
    The proof of Theorem F.1 follows exactly the same steps as the proofs of Theorem 3.4 (for the minimization setting) and Theorem E.4 (for the maximization setting) that were presented in Appendix A and Appendix E respectively. In these proofs, we separately bounded the cost/utility contribution of each MDP by the corresponding surrogate costs/values and simply invoked the feasibility of the optimal adaptive policy at the end. Thus, the exact same proofs imply Theorem F.1 for any feasibility constraint $\mathcal{F}$ and any cost/reward function $h(\cdot)$.
  2. 2.
    The proof of Theorem F.2 follows exactly the same steps as the proofs of Theorem 4.1 (for the minimization setting) and Theorem E.5 that were presented in Appendix B and Appendix E respectively. In particular, none of these proofs used the fact that $\mathcal{F}$ is a matroid constraint at any point, and it is also straightforward to see that they immediately extend for any cost/reward function $h(\cdot)$.

### v2

@@ Appendix I Omitted Proofs from Section 8 (Optional Inspection PB)
We consider the following algorithm, restated from Section 8:
Input : A normalized matroid-max-PBOI instance $\mathbb{I}=(\mathcal{B}_{1},\dots,\mathcal{B}_{n},\mathcal{F})$.
A vector of probabilities $(p_{1},\dots,p_{n})$.
1 Relabel the boxes such that $\mu_{1}\geq\dots\geq\mu_{n}$
$S^{\mathrm{grab}}\leftarrow\{\}$
  set of boxes marked as “grab”
2 For  $i\leftarrow 1,\dots,n$ :
3
        $L_{i}=0$ means we never want to grab box $i$
       Sample $K_{i}\leftarrow\operatorname{Bernoulli}(p_{i})$
        $K_{i}=1$ means provisionally mark box $i$ “grab”
4       If $K_{i}=1\textbf{ and }L_{i}=1$ :
             $S^{\mathrm{grab}}\leftarrow S^{\mathrm{grab}}\cup\{i\}$
              fully mark box $i$ as “grab”
5
6
7
8
9
10Commit to grabbing the boxes $S^{\mathrm{grab}}$ and opening the boxes in $[n]\setminus S^{\mathrm{grab}}$.
11 Run the optimal (index) policy under the resulting commitment.
Algorithm 3 Semilocal Approximation Composition Algorithm
To simplify the notation, we let $W^{\mathrm{open}}_{i}:=W^{*}_{o,i}$ be the surrogate cost of the opening action for box $\mathcal{B}_{i}$. There are two main steps of the proof, each stated and proved in a lemma below. We express both steps in terms of the random variables
$W^{\mathrm{alg}}_{i}=\begin{cases}W^{\mathrm{open}}_{i}&\text{if }K_{i}L_{i}=0\\ \mu_{i}&\text{if }K_{i}L_{i}=1,\end{cases}$
where $K_{i}$ and $L_{i}$ are as defined in Algorithm 3. One can think of $W^{\mathrm{alg}}_{i}$ as the surrogate value of box $i$ conditional on the state of the algorithm at Algorithm 3. The first step, Lemma I.1, is to express the value achieved by Algorithm 3 in terms of $W^{\mathrm{alg}}_{i}$:
Lemma I.1.
[…]
The second step, Lemma I.2, is to compare the resulting expression to an upper bound on the optimal value. This step uses the semilocal $(\alpha,\beta)$-approximation guarantee from Definition 12, which gives us a relationship between $W^{\mathrm{alg}}_{i}$ and $W^{*}_{i}$.
Lemma I.2.
Under the hypotheses of Theorem 8.2,
$\mathbf{E}\bigg{[}\max_{S\in\mathcal{F}}\sum_{i\in S}W^{\mathrm{alg}}_{i}\bigg{]}\geq\mathbf{E}\bigg{[}\max_{S\in\mathcal{F}}\sum_{i\in S}\alpha W^{*}_{i}-\smashoperator[]{\sum_{i\in S^{\mathrm{grab}}}^{}}\beta\mu_{i}\bigg{]},$
Combining the lemmas yields
[…]
  1. From the analogue of Theorem 3.4 for the maximization setting (see Theorem A.4 in Appendix A), we know that the water filling surrogate costs provide an upper bound on the utility of the optimal adaptive policy for any max-CICS instance. Thus:
    $\mathbf{E}\bigg{[}\max_{S\in\mathcal{F}}\sum_{i\in S}W^{*}_{i}\bigg{]}\geq\mathbf{E}[\text{value of optimal policy}].$
  2. Because Algorithm 3 ensures $S^{\mathrm{grab}}\in\mathcal{F}$ by construction, the following policy is feasible: “Compute $S^{\mathrm{grab}}$ as in Algorithm 3, but then simply grab the boxes in $S^{\mathrm{grab}}$.” This algorithm achieves value $\mathbf{E}\big{[}\sum_{i\in S^{\mathrm{grab}}}\mu_{i}\big{]}$, which means
[…]
Proof of Lemma I.1.
  2. If $K_{i}L_{i}=1$ (marked “grab”), $\mathcal{B}^{\prime}_{i}=(\mu_{i},0)$, i.e. the box is free to open and always contains value $\mu_{i}$.
Under instance $\mathbb{I}^{\prime}$, box $i$’s surrogate value is given by $W^{\mathrm{alg}}_{i}$. Since this is an instance of max-CICS over Markov chains, from the counterpart of Theorem 3.2 for maximization (see Theorem A.2 in Appendix A) we have
[…]
The lemma then follows by the law of total expectation.
Proof of Lemma I.2.
We want to prove that
$\mathbf{E}\bigg{[}\max_{S\in\mathcal{F}}\sum_{i\in S}W^{\mathrm{alg}}_{i}\bigg{]}\geq\mathbf{E}\bigg{[}\max_{S\in\mathcal{F}}\sum_{i\in S}\alpha W^{*}_{i}-\smashoperator[]{\sum_{i\in S^{\mathrm{grab}}}^{}}\beta\mu_{i}\bigg{]}.$ (10)
The outline of the proof is as follows. We begin with the left-hand side of 10. Then, for each box $i$, we swap $W^{\mathrm{alg}}_{i}$ with $\alpha W^{*}_{i}$, and also subtract $\beta\mu_{i}$ if box $i$ is marked “grab”. Because each box admits a semilocal $(\alpha,\beta)$-approximation, each of these replacements only decreases the expression’s expected value. After all $n$ replacements, we are left with the right-hand side of 10, as desired. This is the same strategy used by Scully and Doval (2024, Theorem 5.4), but some careful conditioning is needed to account for the $\beta\mu_{i}$ subtractions.
In order to notate the one-by-one replacement outlined above, let
$W^{(j)}_{i}$ =$=\begin{cases}W_{i}^{\mathrm{alg}}&\text{if }i\leq j\\ \alpha W_{i}^{*}&\text{if }i>j,\end{cases}$ $U^{(j)}$ $=\max_{S\in\mathcal{F}}\sum_{i\in S}W^{(j)}_{i}.$
Using this notation and recalling how $S^{\mathrm{grab}}$ is defined in Algorithm 3, we can rewrite our goal 10 as
$\mathbf{E}[U^{(n)}]\geq\mathbf{E}\bigg{[}U^{(0)}-\sum_{i=1}^{n}\beta K_{i}L_{i}\mu_{i}\bigg{]}.$
Therefore, it suffices to show that for all $j\in\{1,\dots,m\}$,
$\mathbf{E}[U^{(j)}]\geq\mathbf{E}\bigg{[}U^{(j-1)}-\beta K_{j}L_{j}\mu_{j}\bigg{]}.$ (11)
We will show 11 using the definition of semilocal approximation (Definition 12). But in order to do so, we need to express each side in terms of a maximum between $W^{\mathrm{alg}}_{j}$ or $\alpha W^{*}_{j}$ and a quantity that is independent of box $j$’s value $X_{j}$ and coin flip $K_{j}$ (this will take the place of the outside option in the definition of semilocal approximation). We express the latter quantity in terms of
[…]
These can both be seen as optimal total surrogate values achievable without box $j$. The difference is that $Y_{\neq j}$ optimizes over sets that exclude $j$, whereas $Z_{\neq j}$ optimizes over sets that include $j$ (but still excludes box $j$’s surrogate value from the sum). With the definitions of $Y_{\neq j}$ and $Z_{\neq j}$ in hand, we can express $U^{(j)}$ and $U^{(j-1)}$ as
So to show 11, it suffices to show
$\mathbf{E}[\max\{W^{\mathrm{alg}}_{j},Y_{\neq j}-Z_{\neq j}\}]\geq\mathbf{E}[\max\{\alpha W^{*}_{j},Y_{\neq j}-Z_{\neq j}\}-\beta K_{j}L_{j}\mu_{j}].$
[…]
The key to showing 12 is observing the following independence facts:
The main obstacle to applying the semilocal approximation condition to 12 is that $W^{\mathrm{alg}}_{j}$ depends on $L_{j}$, which in turn depends on $K_{<j}$. Fortunately, we see from Algorithm 3 that
This is because for all $i$, when executing Algorithm 3, the only randomness the algorithm has used is the past coin flips $K_{<i}$. So to show 12, we split into cases based on whether $L_{j}=0$ or $L_{j}=1$.
[…]
So, by Definition 12,
[…]
$\geq\mathbf{E}[\max\{\alpha W^{*}_{j},y\}]-\beta p_{j}\mu_{j}$
$=\mathbf{E}[\max\{\alpha W^{*}_{j},y\}-\beta K_{j}\mu_{j}]$
[…]
The fact that $(Y_{\neq j},Z_{\neq j})$ is conditionally independent of $(K_{j},X_{j})$ given $K_{<j}$ completes the proof of 12 on the event $L_{j}=1$.
Suppose now that $L_{j}=0$. In this case, we mark box $j$ as “open” regardless of the coin flip $K_{j}$, so instead of using the semilocal approximation condition, we will show that marking box $j$ as “open” does not lose any any potential value. Specifically, we will show the following:
  2. If $L_{j}=0$, then $Y_{\neq j}-Z_{\neq j}\geq\mu_{j}$.
Together with the fact that $\mu_{j}\geq h_{j}$ (which holds by the assumption that all degenerate boxes have been normalized), facts (a) and (b) imply that for any $k_{<j}$ such that $K_{<j}=k_{<j}$ induces $L_{j}=0$ in Algorithm 3,
[…]
which completes the proof of 12 on the event $L_{j}=0$. It remains only to show (a) and (b). Fact (a) holds by definition: recall that $h_{j}$ denotes the smallest value of the outside option $y$ for which the optimal policy in the local game will prefer opening to grabbing. For (b), we will show that if $L_{j}=0$, then $Y_{\neq j}\geq\mu_{j}+Z_{\neq j}$. Let
  2. $S^{\mathrm{grab}}_{<j}=S^{\mathrm{grab}}\cap\{1,\dots,j-1\}$ be the boxes marked “grab” before the $j$-th step in the loop, and
$B^{Y}=(B^{Z}\setminus\{j\})\cup\{k\}.$
But $B^{\mathrm{grab}}\setminus B^{Z}\subseteq S^{\mathrm{grab}}_{<j}$, which means $W^{(j)}_{k}=\mu_{k}\geq\mu_{j}$ since Algorithm 3 iterates over the boxes in decreasing order of mean value. This means
[…]
and the proof is completed.
I.2 Proof of Lemma 8.3
See 8.3
Proof.
To simplify notation we denote $\alpha(\beta)$ as $\alpha$ throughout the proof. By Definition 12, proving this lemma amounts to showing that these $\alpha$ and $\beta$ satisfy:
$(1-p)\mathbf{E}[\max\{W^{*}_{o},y\}]+p\max\{\mu,y\}\geq\mathbf{E}[\max\{\alpha W^{*},y\}]-p\beta\mu$ (13)
for all $y\geq 0$. This is equivalent to showing that
$f(y):=(1-p)\mathbf{E}[\max\{W^{*}_{o},y\}]+p\max\{\mu,y\}-\mathbf{E}[\max\{\alpha W^{*},y\}]+p\beta\mu$
satisfies $f(y)\geq 0$ for all $y\geq 0$. We will first show that it is sufficient to check that $f(0)\geq 0$ and that $f(y)\geq 0$ for $y\geq\mu$. It follows from Definition 11 that at all values $y$ for which $f^{\prime}(y)$ is defined, it is equal to
$(1-p)\operatorname{Pr}\left[y\geq W^{*}_{o}\right]+p\mathbbm{1}(y>\mu)-\mathbbm{1}(y>\alpha h)\operatorname{Pr}\left[y>\alpha W^{*}_{o}\right].$
Observe that
  • if $y<\alpha h<\mu$, then $f^{\prime}(y)\geq 0$,
  • if $\alpha h<y<\mu$, then since $\alpha\leq 1$, we must have $f^{\prime}(y)\leq 0$.
Thus a global minimum of $f(y)$ must be at $y=0$ or on $y\geq\mu$, so it is sufficient to check that 13 is satisfied when $y=0$ and $y\geq\mu$. When $y=0$, 13 reduces to
$(1-p)(\mu-c)+p\mu\geq\alpha\mu-p\beta\mu.$
which holds for the values of $\alpha$ and $p$ given in the lemma. When $y\geq\mu$, 13 reduces to
$(1-p)\mathbf{E}[\max\{W^{*}_{o},y\}]+py\geq\mathbf{E}[\max\{\alpha W^{*},y\}]-p\beta\mu=\alpha\mathbf{E}[\max\{W^{*}_{o},\frac{y}{\alpha}\}]-p\beta\mu$ (14)
where the last equality follows from the fact that $\mathbf{E}[\max\{\alpha W^{*},y\}]=\max\{\mathbf{E}[\max\{\alpha W^{*}_{o},y\}],\alpha\mu\}$. The slope of $\mathbf{E}[\max\{W^{*}_{o},y\}]$ is bounded above by $1$, so
$\alpha\mathbf{E}[\max\{W^{*}_{o},\frac{y}{\alpha}\}]\leq\alpha\mathbf{E}[\max\{W^{*}_{o},y\}]+\alpha\left(\frac{y}{\alpha}-y\right).$
Applying this bound to 14, we find that 14 holds if,
$(\alpha+p-1)(\mathbf{E}[\max\{W^{*}_{o},y\}]-y)\leq p\beta\mu.$ (15)
Since the slope of $\mathbf{E}[\max\{W^{*}_{o},y\}]$ is bounded above by $1$, it must be that $\mathbf{E}[\max\{W^{*}_{o},y\}]-y$ is maximized at $y=0$ where it is equal to $\mu-c$, so 15 holds if,
$(\alpha+p-1)(\mu-c)\leq p\beta\mu.$
Observe that if $p=\alpha\frac{c}{\mu}$, we can rewrite this inequality as
$\alpha\left(1+\frac{c}{\mu}-\frac{c}{\mu-c}\beta\right)\leq 1,$
which is satisfied by the $\alpha$ given in the lemma. ∎